OpenStudy (anonymous):

fint the 8th term of the arithmetic sequence -3.2,-3.4,-3.6,-3.8

OpenStudy (anonymous):

a8 = a+ 7d

OpenStudy (anonymous):

It seems the constant difference is x-.2

OpenStudy (anonymous):

d = - 0.2

OpenStudy (anonymous):

a = -3.2

OpenStudy (anonymous):

a8 = -3.2 + 7* -0.2 = ??

OpenStudy (anonymous):

Find the 9th term of the arithmetic sequence. a10 = 100 and a20 = 50; a9 = _____.